A survey of 1084 Californian males (age 18 and over) indicated more than half did not have a physical exam or blood cholesterol check. A medical researcher plans to sample men in the local community to see if similar results occur. How large a random sample would he need to estimate this proportion to within 0.05 and probability of 0.95?
Answer #1

Solution,

Given that,

=  1 - = 0.5

margin of error = E = 0.05

At 95% confidence level

= 1 - 95%

= 1 - 0.95 =0.05

/2 = 0.025

Z/2 = Z0.025  = 1.96

sample size = n = (Z / 2 / E )2 * * (1 - )

= (1.96 / 0.05)2 * 0.5 * 0.5

= 384.16

sample size = n = 385